# Acceleration of the CASINO quantum Monte Carlo software using graphics processing units and OpenACC

### v1

@@ 1 Introduction
Continuum quantum Monte Carlo (QMC) methods are powerful techniques for finding approximate solutions to the many-body Schrödinger equation. In these methods the quantum many-body problem is tackled directly in the continuous, $3N$-dimensional position basis, with Monte Carlo methods being used to explore and integrate over the high-dimensional configuration space. In this article we focus on the variational and diffusion quantum Monte Carlo (VMC and DMC) methods, which allow one to find the ground state and selected excited states of fermionic and bosonic systems, as implemented in the CASINO software [1]. A major challenge for these methods is the computational expense, which not only constrains the range of systems that can be studied, but also limits the accuracy that can be achieved, as users are forced to use smaller simulation supercells, larger time steps, etc. The high cost of QMC methods is partially offset by the fact that QMC algorithms are inherently parallelizable: one can reduce error bars on estimated quantities by averaging over independent random walks. Hence, provided equilibration time is a negligible fraction of total run time, QMC methods are almost perfect for distributed memory message-passing interface (MPI) parallelization, and CASINO has been shown to scale to tens of thousands of MPI processes [2].
@@
A directives-based approach provides a viable technique for allowing control of offloading from within a single Fortran code base. The OpenACC programming standard and API [5] is supported by Fortran compilers including NVIDIA’s nvfortran, the Cray/HPE Fortran compiler, and GNU’s gfortran. OpenACC is developed and overseen by the OpenACC Organization, a non-profit organization whose members include the NVIDIA Corporation, AMD, HPE, ARM, and Oak Ridge National Laboratory (home to the Frontier supercomputer), as well as many HPC centers and academic institutions. OpenACC compiler directives are used to identify parts of a program that should be compiled for execution on an accelerator device; furthermore these directives are also used to provide much of the functionality specific to programming for GPUs.
@@ 2.2 Ewald summation
and $\Omega_{\rm s}$ is the volume of the simulation cell, $\{{\bf R}_{\rm s}\}$ are the simulation-cell lattice points, $\{{\bf G}_{\rm s}\}$ are the simulation-cell reciprocal lattice points, and $\gamma$ is a constant that can be chosen to minimize the cost of evaluating Eq. (3) [22]. Similar expressions can be derived for Ewald interactions in two dimensions [23, 24, 25].
It is clear that the cost of evaluating the Ewald interaction in Eq. (3) is vastly greater than the cost of evaluating $1/r$. Equation (3) is suitable for offloading to GPUs because the evaluation of the series is dominated by floating point operations and both the real- and reciprocal-space series are inherently parallelizable; moreover, much of the data required can be moved to the GPU once, at the outset of the calculation. Offloading of Eq. (3) is investigated in Section 5.1.
@@ 2.3 Long-range Jastrow terms
[…]
@@
so only three complex exponentials are required to compute all the required cosines for Eq. (7). The powers of these exponentials are computed and buffered. The required number of reciprocal lattice points is typically $\sim$10$\sim$100. We note that for large plane-wave expansions there is scope to reduce the computational cost of the $p$ term by means of a blip (B-spline) re-representation.
@@ 6.2 Accuracy of SP arithmetic in QMC calculations
Switching from DP to SP makes surprisingly little difference to VMC and DMC results when a blip basis set is used for the orbitals. For example, the VMC total energy of a $2\times 2\times 2$ supercell of bulk Si at the simulation-cell Baldereschi point using Trail-Needs Si pseudopotentials was found to be $-7.875606(14)$ Ha per primitive cell with SP arithmetic and $-7.875615(10)$ Ha per primitive cell with DP arithmetic. In these calculations, a Slater-Jastrow trial wave function was used, with $u$, $\chi$, $f$, and $p$ terms in the Jastrow exponent and complex Perdew-Burke-Ernzerhof orbitals represented in a blip basis set in the Slater determinant. The parameters in the Jastrow factor were optimized by energy minimization with DP arithmetic. The corresponding DMC energies at a time step of 0.04 Ha-1 were found to be $-7.88720(4)$ and $-7.88713(4)$ Ha per primitive cell using SP and DP arithmetic, respectively. Similarly excellent agreement between SP and DP QMC results is found in a $2\times 2\times 2$ supercell of Si at the simulation-cell $\Gamma$ point, where the trial wave function is real. These results demonstrate that SP arithmetic is perfectly adequate for at least some QMC calculations.
@@
On the other hand, when an optimized Gaussian basis set is used for a $2\times 2\times 2$ supercell of Si at the $\Gamma$ point with a Troullier-Martins Si pseudopotential, the Slater-Jastrow VMC SP and DP energies are $-7.88573(3)$ and $-7.88283(3)$ Ha per primitive cell, respectively. These differ significantly, indicating that Gaussian basis sets are more challenging for SP arithmetic. It was verified that this difference is not affected by reasonable changes to the tolerance controlling the distances beyond which Gaussian basis functions are truncated to zero. The difference in SP and DP total energies is equal to the corresponding difference in kinetic energies, while the potential energies are in statistical agreement, implying that the evaluation of the Laplacians of orbitals expanded in a Gaussian basis set is the part that is particularly sensitive to the choice of precision.
@@
Furthermore, even when a blip basis set is used, wave-function optimization is less effective with SP arithmetic than with DP arithmetic. Wave function optimization is usually computationally cheaper than DMC, but is the most technically challenging aspect of QMC work and the most costly in terms of human time. Re-optimization by energy minimization with SP arithmetic does not significantly alter or spoil the previously optimized trial wave function used in the results quoted above for a $2\times 2\times 2$ supercell of Si at the Baldereschi point. However, optimization of the parameters in that trial wave function from scratch with SP arithmetic yields a VMC energy of $-7.86165(9)$ Ha per primitive cell, which is significantly higher than the VMC energy of $-7.875606(14)$ Ha per primitive cell obtained using energy minimization with DP arithmetic.
@@ 6.3 Speedups due to offloading with SP arithmetic
Activity Run type $N$ Runtime (s) Speedup vs. CPU
@@ 9 Conclusions
We have investigated the speedup to the CASINO code that can result from using OpenACC to offload various aspects of a quantum Monte Carlo calculation with a Slater-Jastrow trial wave function to a GPU. For electron gases in two and three dimensions we find that, on a POWER9 CPU with an NVIDIA V100 GPU, offloading the evaluation of Ewald interactions is beneficial when more than about 350 electrons are present; with 1000 electrons, offloading accelerates the evaluation of Ewald interactions by a factor of $\sim 2.5$. Likewise, offloading the computation of long-range two-body terms in the Jastrow correlation factor is beneficial with more than about 150 electrons, speeding up Jastrow factor evaluation by a factor of (again) $\sim 2.5$ with 1000 electrons. The use of SP arithmetic increases the speedup due to offloading to a factor of about three (on top of the speedup of about 1.3 to the CPU version due to the use of SP). In electron gases, the evaluation of two-body Jastrow terms and Ewald interactions account for a significant fraction of the overall runtime, and so these accelerations translate into a similar speedup for the entire calculation.
@@
These results were obtained by comparing timings for a single process with a single thread running on a single CPU, with and without offloading to a single GPU. An important issue to investigate is the optimal use of GPUs on compute nodes with multiple MPI processes competing for the same set of GPUs. Furthermore it would be desirable to investigate the benefits of offloading on a range of different hardware. Finally, this work has focussed on the kinds of QMC problem of interest in condensed-matter physics: we have studied electron gas models and a 2D semiconductor. In large molecular systems, electron-electron-nucleus ($f$) Jastrow terms are relatively important, and an obvious future development would be to investigate the offloading of these Jastrow terms.

### v2

@@ 1 Introduction
Continuum quantum Monte Carlo (QMC) methods are powerful techniques for finding approximate solutions to the many-body Schrödinger equation. In these methods the quantum many-body problem is tackled directly in the continuous, $3N$-dimensional position basis, with Monte Carlo methods being used to explore and integrate over the high-dimensional configuration space. In this article we focus on the variational and diffusion quantum Monte Carlo (VMC and DMC) methods, which allow one to find the ground state and selected excited states of fermionic and bosonic systems, as implemented in the CASINO software [1, 2]. A major challenge for these methods is the computational expense, which not only constrains the range of systems that can be studied, but also limits the accuracy that can be achieved, as users are forced to use smaller simulation supercells, larger time steps, etc. The high cost of QMC methods is partially offset by the fact that QMC algorithms are inherently parallelizable: one can reduce error bars on estimated quantities by averaging over independent random walks. Hence, provided equilibration time is a negligible fraction of total run time, QMC methods are almost perfect for distributed memory message-passing interface (MPI) parallelization, and CASINO has been shown to scale to tens of thousands of MPI processes [3].
@@
A directives-based approach provides a viable technique for allowing control of offloading from within a single Fortran code base. The OpenACC programming standard and API [6] is supported by Fortran compilers including NVIDIA’s nvfortran, the Cray/HPE Fortran compiler, and GNU’s gfortran. OpenACC is developed and overseen by the OpenACC Organization, a non-profit organization whose members include the NVIDIA Corporation, AMD, HPE, ARM, and Oak Ridge National Laboratory (home to the Frontier supercomputer), as well as many high-performance computing centers and academic institutions. OpenACC fully supports Fortran, C, and C++. OpenACC compiler directives are used to identify parts of a program that should be compiled for execution on an accelerator device; furthermore these directives are also used to provide much of the functionality specific to programming for GPUs.
@@ 2.2 Ewald summation
and $\Omega_{\rm s}$ is the volume of the simulation cell, $\{{\bf R}_{\rm s}\}$ are the simulation-cell lattice points, $\{{\bf G}_{\rm s}\}$ are the simulation-cell reciprocal lattice points, and $\gamma$ is a constant that can be chosen to minimize the cost of evaluating Eq. (3) [23]. Similar expressions can be derived for Ewald interactions in two dimensions [24, 25, 26]:
$v_{\rm E}({\bf r})$ $=$ $\frac{\pi}{A_{\rm s}}\sum_{{\bf G}_{\rm s}\neq{\bf 0}}\frac{e^{zG_{\rm s}}\operatorname{erfc}\left(\frac{G_{\rm s}}{2\sqrt{\gamma}}+z\sqrt{\gamma}\right)+e^{-zG_{\rm s}}\operatorname{erfc}\left(\frac{G_{\rm s}}{2\sqrt{\gamma}}-z\sqrt{\gamma}\right)}{G_{\rm s}}e^{i{\bf G}_{\rm s}\cdot{\bf r}}$ (5)
${}+\sum_{{\bf R}_{\rm s}}\frac{\operatorname{erfc}\big(\sqrt{\gamma}|{\bf r}-{\bf R}_{\rm s}|\big)}{|{\bf r}-{\bf R}_{\rm s}|}-\frac{2\pi}{A_{\rm s}}\left[z\operatorname{erf}(z\sqrt{\gamma})+\frac{e^{-\gamma z^{2}}}{{(\gamma\pi)}^{\frac{1}{2}}}\right]$
$v_{\rm M}$ $=$ $\frac{2\pi}{A_{\rm s}}\sum_{{\bf G}_{\rm s}\neq{\bf 0}}\frac{\operatorname{erfc}\left(\frac{G_{\rm s}}{2\sqrt{\gamma}}\right)}{G_{\rm s}}+\sum_{{\bf R}_{\rm s}\neq{\bf 0}}\frac{\operatorname{erfc}\left(\sqrt{\gamma}R_{\rm s}\right)}{R_{\rm s}}-\frac{2\sqrt{\gamma}}{\pi^{\frac{1}{2}}}-\frac{2\pi^{\frac{1}{2}}}{A_{\rm s}\sqrt{\gamma}},$ (6)
where $A_{\rm s}$ is the area of the simulation cell.
It is clear that the cost of evaluating the Ewald interaction in Eqs. (3) or (5) is vastly greater than the cost of evaluating $1/r$. Equations (3) and (5) are suitable for offloading to GPUs because the evaluation of the series is dominated by floating point operations and both the real- and reciprocal-space series are inherently parallelizable; moreover, much of the data required can be moved to the GPU once, at the outset of the calculation. Offloading of Ewald interactions is investigated in Section 5.1.
@@ 2.3 Long-range Jastrow terms
where $\{\alpha_{i}\}$ are optimizable parameters for $i=0$ and $i=2,\ldots,N_{u}$, while $\alpha_{1}$ is determined by the Kato cusp conditions, $L_{u}$ is a cutoff length that must be less than half the distance from the origin to the closest nonzero simulation-cell lattice point, $\Theta$ is the Heaviside function and $C$ is the number of derivatives of $u(r)$ that are well-defined at $r=L_{u}$; usually we have $C=2$ or $C=3$. Typically the expansion order $N_{u}$ is chosen to be in the range 8–12. The $\{a_{m}\}$ are optimizable coefficients for each star $\bigstar_{m}$ of simulation-cell reciprocal-lattice points ${\bf G}_{\rm s}$; the “+” superscript indicates that only one ${\bf G}_{\rm s}$ vector out of each $\pm{\bf G}_{\rm s}$ pair is included in the sum. Typically $N_{p}=3$1515 stars of ${\bf G}_{\rm s}$ vectors are used.
@@
so only three complex exponentials are required to compute all the required cosines for Eq. (9). The powers of these exponentials are computed and buffered. The required number of reciprocal lattice points is typically $\sim$$\sim10$0. We note that for large plane-wave expansions there is scope to reduce the computational cost of the $p$ term by means of a blip (B-spline) re-representation.
@@ 6.2 Accuracy of SP arithmetic in QMC calculations
Switching from DP to SP makes surprisingly little difference to VMC and DMC results when a blip basis set is used for the orbitals. For example, the VMC total energy of a $2\times 2\times 2$ supercell of bulk Si at the simulation-cell Baldereschi point using Trail-Needs Si pseudopotentials was found to be 7.875606(14)-7.875606(14) Ha per primitive cell with SP arithmetic and 7.875615(10)-7.875615(10) Ha per primitive cell with DP arithmetic. In these calculations, a Slater-Jastrow trial wave function was used, with $u$, $\chi$, $f$, and $p$ terms in the Jastrow exponent and complex Perdew-Burke-Ernzerhof orbitals represented in a blip basis set in the Slater determinant. The parameters in the Jastrow factor were optimized by energy minimization with DP arithmetic. The corresponding DMC energies at a time step of 0.04 Ha-1 were found to be 7.88720(4)-7.88720(4) and 7.88713(4)-7.88713(4) Ha per primitive cell using SP and DP arithmetic, respectively. Similarly excellent agreement between SP and DP QMC results is found in a $2\times 2\times 2$ supercell of Si at the simulation-cell $\Gamma$ point, where the trial wave function is real. These results demonstrate that SP arithmetic is perfectly adequate for at least some QMC calculations.
@@
On the other hand, when an optimized Gaussian basis set is used for a $2\times 2\times 2$ supercell of Si at the $\Gamma$ point with a Troullier-Martins Si pseudopotential, the Slater-Jastrow VMC SP and DP energies are 7.88573(3)-7.88573(3) and 7.88283(3)-7.88283(3) Ha per primitive cell, respectively. These differ significantly, indicating that Gaussian basis sets are more challenging for SP arithmetic. It was verified that this difference is not affected by reasonable changes to the tolerance controlling the distances beyond which Gaussian basis functions are truncated to zero. The difference in SP and DP total energies is equal to the corresponding difference in kinetic energies, while the potential energies are in statistical agreement, implying that the evaluation of the Laplacians of orbitals expanded in a Gaussian basis set is the part that is particularly sensitive to the choice of precision.
@@
Furthermore, even when a blip basis set is used, wave-function optimization is less effective with SP arithmetic than with DP arithmetic. Wave function optimization is usually computationally cheaper than DMC, but is the most technically challenging aspect of QMC work and the most costly in terms of human time. Re-optimization by energy minimization with SP arithmetic does not significantly alter or spoil the previously optimized trial wave function used in the results quoted above for a $2\times 2\times 2$ supercell of Si at the Baldereschi point. However, optimization of the parameters in that trial wave function from scratch with SP arithmetic yields a VMC energy of 7.86165(9)-7.86165(9) Ha per primitive cell, which is significantly higher than the VMC energy of 7.875606(14)-7.875606(14) Ha per primitive cell obtained using energy minimization with DP arithmetic.
@@ 6.3 Speedups due to offloading with SP arithmetic
Activity Run type NN Runtime (s) Speedup vs. CPU
@@ 9 Conclusions
We have investigated the speedup to the CASINO code that can result from using OpenACC to offload various aspects of a quantum Monte Carlo calculation with a Slater-Jastrow trial wave function to a GPU. For electron gases in two and three dimensions we find that, on a POWER9 CPU with an NVIDIA V100 GPU, offloading the evaluation of Ewald interactions is beneficial when more than about 350 electrons are present; with 1000 electrons, offloading accelerates the evaluation of Ewald interactions by a factor of $\sim 2.5$. Likewise, offloading the computation of long-range two-body terms in the Jastrow correlation factor is beneficial with more than about 150 electrons, speeding up Jastrow factor evaluation by a factor of (again) $\sim 2.5$ with 1000 electrons. The use of SP arithmetic increases the speedup due to offloading to a factor of about three (on top of the speedup of about 1.3 to the CPU version due to the use of SP). In electron gases, the evaluation of two-body Jastrow terms and Ewald interactions account for a significant fraction of the overall runtime, and so these accelerations translate into a similar speedup for the entire calculation. A speedup of $2.5$ reduces random errors on a QMC calculation by a factor of $1/\sqrt{2.5}\approx 0.63$ for a given computational resource. Or, for a given error bar on the total energy and a given computational resource, CASINO could be used to study a system that is $2.5^{1/3}\approx 1.4$ times larger.
@@
These results were obtained by comparing timings for a single process with a single thread running on a single CPU, with and without offloading to a single GPU. Our approach of offloading two-body Jastrow terms is applicable in VMC, DMC, or reptation quantum Monte Carlo software, while offloading of Ewald interactions is also applicable in path integral Monte Carlo calculations. An important issue to investigate is the optimal use of GPUs on compute nodes with multiple MPI processes competing for the same set of GPUs. Furthermore it would be desirable to investigate the benefits of offloading on a range of different hardware. Finally, this work has focussed on the kinds of QMC problem of interest in condensed-matter physics: we have studied electron gas models and a 2D semiconductor. In large molecular systems, electron-electron-nucleus ($f$) Jastrow terms are relatively important, and an obvious future development would be to investigate the offloading of these Jastrow terms.